You are taking a multiple-choice test that has 12 questions. Each of the questions has 5 choices, with one correct choice per question. If you select one of these options per question and leave nothing blank, in how many ways can you answer the questions?
The answer is 244,140,625 but why?

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

You are taking a multiple-choice test that has 12 questions. Each of the questions has 5 choices, with one correct choice per question. If you select one of these options per question and leave nothing blank, in how many ways can you answer the questions?
---
Each question has 5 possible choices for an answer>
1st question: 5 ways
2nd question: 5 ways
.......
12th question: 5 ways
----
Total # of possible answers 5*5*5...*5*5 = 5^12 = 244,140,625
-----
Cheers,
Stan H.
=========================
The answer is 244,140,625 but why?

Answer by edjones(8007)

5^12=244140625
.
Ed